### v3

# Compact Kähler manifolds with quasi-positive holomorphic sectional curvature

Abstract. In this paper, we establish a “pseudo-effective” version of the holonomy principle for compact Kähler manifolds with nonnegative holomorphic sectional curvature. As applications, we prove that if a compact complex manifold $M$ admits a Kähler metric with nonnegative holomorphic sectional curvature and $(M,\omega)$ has no nonzero truly flat tangent vector at some point (which is satisfied when the holomorphic sectional curvature is quasi-positive), $M$ is projective and rationally connected. This answers a problem raised by Yang and Matsumura and extends Yau’s conjecture. We also get the structure theorem of a compact simply connected Kähler manifold with nonnegative holomorphic sectional curvature. Furthermore, we show that a compact Kähler manifold admits a Hermitian metric with positive real bisectional curvature is a projective and rationally connected manifold.
AMS Mathematics Subject Classification. 53C07, 58E15
Keywords and phrases. Kähler metric, holomorphic sectional curvature, projective, rational connectedness
1. Introduction
A projective manifold $M$ is called rationally connected if any two points on $M$ can be connected by some rational curve. The rational connectedness is an important and fundamental concept in algebraic geometry, and many people have given the criteria for it ([8, 17, 31, 23, 7, 22]). It is also interesting to give geometric interpretations for rational connectedness. Let $\omega$ be a Kähler metric on $M$, the holomorphic sectional curvature of $\omega$ along the nonzero $V\in T^{1,0}M$ is defined by
$H(V)=\frac{R(V,\overline{V},V,\overline{V})}{|V|_{\omega}^{4}},$
where $R$ is the Riemannian curvature tensor of $\omega$. S.-T. Yau proposed the following well-known conjecture in his problem section.
Conjecture 1.1 ([39, Problem 47]).
If a compact complex manifold $M$ admits a Kähler metric $\omega$ with positive holomorphic sectional curvature, then $M$ is projective and rationally connected.
By studying MRC fibration, Heier and Wong ([18]) confirmed Yau’s conjecture for projective manifolds with quasi-positive holomorphic sectional curvature. Then Yang ([36, 37]) proved that a Kähler manifold $M$ with positive holomorphic sectional curvature satisfies $H^{p,0}_{\bar{\partial}}(M)=H^{0,2}_{\bar{\partial}}(M)=0$ by introducing the notion of RC-positivity and affirmly confirmed Yau’s conjecture, but it’s seems difficult to apply his method to the case of quasi-positive holomorphic sectional curvature. There are some other interesting curvature conditions toward rational connectedness. L.Ni ([28]) introduced the concept of $k$-Ricci curvature for compact Kähler manifold and generalized Yang’s method to show that a compact Kähler manifold with positive $k$-Ricci curvature is projective and rationally connected. Recently, Li, Zhang and the second author ([24, Theorem 1.14]) established the equivalence between mean curvature (quasi-)positivity and rational connectedness for compact Kähler manifolds.
In [25, 26], Matsumura established the structure theorem for a projective manifold $X$ with nonnegative holomorphic sectional curvature by studying the structure of MRC fibration: the universal cover of $X$ splits isometrically as $F\times\mathbb{C}^{n-k}$ for some rationally connected manifold $F$. As a corollary, he generalized Yau’s conjecture for projective manifolds ([26, Theorem 1.2]): if a projective manifold $M$ admits a Kähler metric $\omega$ with nonnegative holomorphic sectional curvature and $(M,\omega)$ has no nonzero truly flat tangent vector at some point (which is satisfied when the holomorphic sectional curvature is quasi-positive), then $M$ is rationally connected. Following the definition in [20], a tangent vector $V\in T_{x_{0}}^{1,0}M$ is said to be truly flat at $x_{0}$ with respect to $\omega$ if
$R(V,\overline{X},Y,\overline{Z})=0$
for all $X,Y,Z\in T_{x_{0}}^{1,0}M$. Motivated by these results, Matsumura and Yang proposed the following natural generalization.
Conjecture 1.2 ([37, Conjecture 1.9], [27, Problem 3.3]).
Let $M$ be a compact Kähler manifold. If $M$ admits a Kähler metric (or more generally a Hermitian metric) with nonnegative holomorphic sectional curvature and there has no nonzero truly flat tangent vector at some point (or the holomorphic sectional curvature is quasi-positive), then $M$ is projective and rationally connected.
More generally, Matsumura proposed the following extension and suggested to investigate Albanese maps instead of MRC fibrations for compact Kähler manifolds. We refer the readers to [27, Section 3], [26, Section 5] for more detailed discussion.
Problem 1.3 ([27, Problem 3.5]).
Can the structure theorem for projective manifolds with nonnegative holomorphic sectional curvature be generalized to compact Kähler manifolds?
In this paper, we establish the following “pseudo-effective” version of the holonomy principle for nonnegative holomorphic sectional curvature (cf. [8, Theorem 1.5]), which is the first main result of our paper.
Theorem 1.4.
Let $(M,\omega)$ be a compact Kähler manifold with nonnegative holomorphic sectional curvature. Assume that there exists an invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ that is pseudo-effective as a line bundle, then the following statements holds:
  • (1)
    There exists some nonzero truly flat tangent vectors at any point of $M$ determined by $\mathcal{F}$.
  • (2)
    $\mathcal{F}$ is flat as a line bundle and is invariant under parallel transport by the connection $D$ of $\Lambda^{p,0}M$ induced by Chern connection of $(T^{1,0}M,\omega)$; in fact, the local holonomy $H$ of $(T^{1,0}M,\omega)$ acts trivally on $\mathcal{F}$.
The proof of Theorem 1.4 is based on an integral inequality associated with $\mathcal{F}$ (Proposition 3.1). As a direct application of Theorem 1.4 (1) and Campana-Demailly-Peternell’s criterion for rational connectedness, we get the second main result of our paper, which confirms Conjecture 1.2 for the Kähler case and extends Yau’s conjecture to the quasi-positive case. Our approach is independent of those of Heier-Wong ([18]) and Matsumura ([26]) for projective manifolds.
Theorem 1.5.
Let $(M,\omega)$ be a compact Kähler manifolds. If the holomorphic sectional curvature holomorphic sectional curvature is nonnegative and $(M,\omega)$ has no nonzero truly flat tangent vector at some point, then $M$ is projective and rationally connected.
Remark 1.6.
Statement (2) of Theorem 1.4 is same as the statement of [8, Theorem 1.5] for nonnegative Ricci curvature, but the proof is more involved. Based on this statement, Campana-Demailly-Peternell’s criterion for rational connectedness (Lemma 2.5), Cheeger-Gromoll’s splitting theorem ([9]), De Rham’s splitting theorem ([15]) and Berger’s classification ([2]) for irreducible holonomy, Campana, Demailly and Peternell established the structure theorem of a compact Kähler manifold $M$ with nonnegative Ricci curvature ([8, Theorem 1.4]): the universal cover $\widetilde{M}$ of $M$ can splits isometrically as $\mathbb{C}^{q}\times\prod Y_{j}\times\prod S_{k}\times\prod Z_{l}$, where $Y_{j}$, $S_{k}$ and $Z_{l}$ are compact simply connected Kähler manifolds of respective dimensions $n_{j}$, $n^{\prime}_{k}$, $n^{\prime\prime}_{l}$ with irreducible holonomy, $Y_{j}$ being Calabi-Yau manifolds (holonomy $SU(n_{j})$), $S_{k}$ holomorphic symplectic manifolds (holonomy $Sp(n^{\prime}_{k}/2)$), and $Z_{l}$ rationally connected manifolds with $K_{Z_{l}}^{-1}$ semipositive (holonomy $U(n^{\prime\prime}_{l})$, unless $Z_{l}$ is Hermitian symmetric of compact type).
Statement (1) of Theorem 1.4 ensures that factors of irreducible holonomy $H=SU(n_{j})$ or $Sp(\frac{n_{k}^{\prime}}{2})$ can be excluded for a compact Kähler manifold with nonnegative holomorphic sectional curvature.
Theorem 1.7.
Let $(M,\omega)$ be a compact simply connected Kähler manifold with nonnegative holomorphic sectional curvature. Then $M$ admits a holomorphic and isometric splitting $(M,\omega)\simeq\prod(M_{l},\omega_{l})$ where $M_{l}$ are rationally connected manifolds with nonnegative holomorphic sectional curvature and irreducible holonomy (holonomy $U(n_{l})$ unless $M_{l}$ is Hermitian symmetric of compact type).
If the following splitting theorem for nonnegative holomorphic sectional curvature established, Problem 1.3 will be resolved.
Question 1.8.
Let $(M,\omega)$ be a compact Kähler manifold with nonnegative holomorphic sectional curvature. Can the universal cover $\widetilde{M}$ of $M$ splits isometrically as $N\times\mathbb{C}^{k}$ for some compact Kähler manifold $N$?
For the Hermitian category, Yang ([36, Corollary 1.10]) proved that a compact Kähler surface admits a Hermitian metric with positive holomorphic sectional curvature is projective and rationally connected. In [38], Yang-Zheng introduced the concept of real bisectional curvature as in the following way and generalized Wu-Yau’s Theorem ([19, 34, 35, 16]) to the Hermitian case. The underlying reason for the extension is that the Schwarz lemma works the same when the target metric is only Hermitian but has nonpositive real bisectional curvature on a Kähler background. Also, they proposed the following generalization of Yau’s Conjecture 1.1 for the Hermitian case ([38, Conjecture 1.6 (b)]).
Conjecture 1.9.
Let $(M,h)$ be a Hermitian manifold with positive real bisectional curvature. Then M is a projective and rationally connected manifold.
We say that a Hermitian manifold $(M,h)$ has positive real bisectional curvature denoted by $B_{h}>0$ if for any $x\in M$, any unitary frame $e=\{e_{1},\cdots,e_{n}\}$ at $x$ and any nonnegative constants $a=\{a_{1},\cdots,a_{n}\}$ with $|a|^{2}=a_{1}^{2}+\cdots+a_{n}^{2}>0$, it holds that
$B_{h}(e,a)=\frac{1}{|a|^{2}}\sum\limits_{i,j=1}^{n}R_{i\bar{i}j\bar{j}}a_{i}a_{j}>0,$
where $R$ is the chern curvature tensor of $h$. The term $B_{h}\geq 0,\ B_{h}<0$ or $B_{h}\leq 0$ are defined similarly. This curvature condition is the same as positive holomorphic sectional curvature when the metric $h$ is Kähler and slightly stronger in the non-Kähler case (cf. [38]). Under the entra assumption that $M$ is Kählerian, Conjcture 1.9 is just a weaker version of Conjecture 1.2. In this special case, Tang ([33, Theorem 1.4]) confirmed the Conjecture 1.9 for $\dim M=3$ and proved that $M$ is projective for $\dim M\geq 1$. Since the curvature term of Bochner-type formula in the proof of Proposition 3.1 can be controlled by real bisectional curvature, we confirm Conjecture 1.9 under the additional assumption that $M$ is Kählerian.
Theorem 1.10.
Let $(M,h)$ be a compact Hermitian manifold with positive real bisectional curvature. Then for any holomorphic vector bundle $E$ on $M$, there exists a constant $C_{E}>0$ such that
$H^{0}(M,((T^{1,0}M)^{*})^{\otimes p}\otimes E^{\otimes l})=0\ \ \text{for all $p,l\in\mathbb{N}_{+}$ with $p\geq C_{E}l$.}$
In particular, if $M$ is Kählerian, $M$ is a projective and rationally connected manifold.
This paper is organized as follows. In section 2, we recall some basic definitions and facts. In section 3, we give a proof of Theorem 1.4 by establishing an integral inequality. In section 4 , we give a proof of Theorem 1.5 and Theorem 1.7. In section 5, we give a proof of Theorem 1.10.
2. Preliminaries
For reader’s convenience, we gather some basic definitions and useful facts known to experts, which will be referenced later in the text.
2.1. Pseudo-effective line bundle and plurisubharmonic function
We first introduce the concept of pseudo-effective line bundle (see e.g. [11, 14]).
Definition 2.1.
A singular Hermitian metric $h$ on a line bundle $L$ is a metric which is given in any trivialization $\theta:F|_{\Omega}\rightarrow\Omega\times\mathbb{C}$ by
$\|\xi\|_{h}=|\theta(\xi)|e^{-\varphi(x)},\ x\in\Omega,\xi\in L_{x}$
where $\varphi\in L^{1}_{loc}(\Omega)$, called the weight of the metric with respect to the trivialization $\theta$.
Definition 2.2.
Let $L$ be a line bundle on a compact complex manifold $X$. We say that $L$ is pseudo-effective if $L$ can be equipped with a singular Hermitian metric $h$ with $T=\frac{\sqrt{-1}}{2\pi}\Theta_{L,h}=\sqrt{-1}\partial\bar{\partial}\varphi\geq 0$ as a current, equivalently, the weight $\varphi\in Psh(\Omega)\cap L_{loc}^{1}(\Omega)$ in any trivalization.
Let $\Omega$ be an open subset of $\mathbb{C}^{n}$. We have the following approxiamation properties for plurisubharmonic functions (see e.g. [13, Chapter 1 and Chapter 3]).
Lemma 2.3.
Let $u\in Psh(\Omega)\cap L^{1}_{loc}(\Omega)$. If $(\rho_{\epsilon})$ is a family of smoothing kernels, then $u*\rho_{\epsilon}\in Psh(\Omega_{\epsilon})\cap C^{\infty}(\Omega_{\epsilon})$, the family $(u*\rho_{\epsilon})$ is decreasing as $\epsilon\rightarrow 0$ and $\lim_{\epsilon\rightarrow 0}u*\rho_{\epsilon}=u,$ where $\Omega_{\epsilon}=\{x\in\Omega:d(x,\partial\Omega)>\epsilon\}$.
Lemma 2.4.
([3, Theorem 2.1]) Let $\{v_{j}^{1}\},\cdots,\{v_{j}^{k}\}$ be a decreasing sequences of functions in $Psh(\Omega)\cap L^{\infty}_{loc}(\Omega)$ and assume that for all $z\in\Omega$,
$\lim\limits_{j\rightarrow\infty}v_{j}^{i}=v^{i}\in Psh(\Omega)\cap L^{\infty}_{loc}(\Omega),\ 1\leq i\leq k.$
Then
$\lim\limits_{j\rightarrow\infty}v_{j}^{1}dd^{c}v_{j}^{2}\wedge\cdots\wedge dd^{c}v_{j}^{k}=v^{1}dd^{c}v^{2}\wedge\cdots dd^{c}v^{k},$
where the limits is in the weak topology on the currents of bigdegree $(k,k)$ on of order zero.
2.2. Some criterions for rational connectedness
Based on the uniruledness criterion estabished in ([6]), Campana-Demailly-Peternell establish the following criterions for rational connectedness.
Lemma 2.5 ([8, Theorem 1.1]).
Let $M$ be a projective manifold. The following properties are equivalent.
  1. (a)
    $M$ is rationally connected.
  2. (b)
    For every invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$, $1\leq p\leq n$, $\mathcal{F}$ is not pseudo-effective.
  3. (c)
    For every invertible subsheaf $\mathcal{F}\subset\mathcal{O}(T^{1,0}M^{*})^{\otimes p})$, $p\geq 1$, $\mathcal{F}$ is not pseudo-effective.
  4. (d)
    For any ample line bundle $A$ on $X$, there exists a constant $C_{A}>0$ such that
    $H^{0}(M,(T^{1,0}M^{*})^{\otimes m}\otimes A^{\otimes k})=0$
    for all $m,k\in\mathbb{N}^{*}$ with $m\geq C_{A}k$.
2.3. Some properties of curvature tensor
Let us recall some algebraic properties about curvature tensor.
Lemma 2.6.
(Royden’s observation, [32, P552]) Let $\xi_{1},\cdots,\xi_{k}$ be $k$ orthonormal tangent vectors. If $S(\xi,\bar{\eta},\zeta,\bar{\omega})$ is a symmetric bihermitian form on some complex vector space $V$ (i.e., $S(\xi,\bar{\eta},\zeta,\omega)=S(\zeta,\bar{\eta},\xi,\bar{\omega})$ and $S(\xi,\bar{\omega},\zeta,\bar{\eta})$), such that for all $\xi\in V$, we have
$S(\xi,\bar{\xi},\xi,\bar{\xi})\leq K|\xi|^{4},$
for some constant $K$, then
$\sum\limits_{\alpha,\beta}S(\xi_{\alpha},\bar{\xi}_{\alpha},\xi_{\beta},\bar{\xi}_{\beta})\leq\frac{1}{2}K[\sum\limits|\xi_{\alpha}|^{4}+(\sum\limits|\xi_{\alpha}|^{2})^{2}].$
Lemma 2.7.
([37, Lemma 1.6]) Let $(M,\omega)$ be a compact Kähler manifold and $x_{0}$ be an arbitrary point on $M$. Let $e_{1}\in T_{x_{0}}^{1,0}M$ be a unit vector which minimizes the holomorphic sectional curvature $H$ of $\omega$ at $x_{0}$, then
$2R(e_{1},\overline{e_{1}},W,\overline{W})\geq(1+|\langle W,\overline{e_{1}}\rangle_{g}|^{2})R(e_{1},\overline{e_{1}},e_{1},\overline{e_{1}})$ (2.1)
for any unit vector $W\in T_{x_{0}}^{1,0}M$.
Lemma 2.8.
([20, Lemma 2.1]) Let $(M,\omega)$ be a compact Kähler manifold and $x_{0}$ be an arbitrary point on $M$. Assume that the holomorphic sectional curvature $H$ of $\omega$ is nonnegative at $x_{0}$. If a nonzero tangent vector $V\in T^{1,0}_{x_{0}}M$ such that $R(V,\overline{V},W,\overline{W})=0$ for any $W\in T^{1,0}_{x_{0}}(M)$, then
$R(V,\overline{X},Y,\overline{Z})=0$ (2.2)
for any $X,Y,Z\in T_{x_{0}}^{1,0}M$, i.e. $V$ is a nonzero truly flat tangent vector at $x_{0}$.
3. Proof of Theorem 1.4
In this subsection, we prove Theorem 1.4. The proof is divided into two parts: establishing an integral inequality and performing a pointwise computation. The method is based on integration by parts, Bochner-type formula and algebraic properties of holomorphic sectional curvature.
Setting.
Let $(M,\omega)$ be a compact Kähler manifold. Assume that $\mathcal{F}$ is an invertible sheaf of $\mathcal{O}(\Lambda^{p,0}M)$ which is pseudo-effective as a line bundle. Then there exist a finite covering $\{U_{\alpha}\}$ consists of coordinate neighborhoods and local holomorphic $(p,0)$-forms $\eta_{\alpha}$ generating $\mathcal{F}|_{U_{\alpha}}$, $\mathcal{F}$ is associated with the Čech cocycle $b_{\alpha\beta}$ in $\mathcal{O}_{M}^{*}(U_{\alpha})$ such that $\eta_{\beta}=b_{\alpha\beta}\eta_{\alpha}$ and there exists a singular Hermitian metric $h=e^{-\varphi}$ of $\mathcal{F}$ defined by a collection of plurisubharmonic functions $\varphi_{\alpha}\in\mathrm{PSH}(U_{\alpha})\cap L_{loc}^{1}(U_{\alpha})$ such that $e^{-\varphi_{\beta}}=|b_{\alpha\beta}|^{2}e^{-\varphi_{\alpha}}$. We can define a globally bounded measurable function $\psi=e^{\varphi_{\alpha}}|\eta_{\alpha}|^{2}_{g}$ as in ([8]), which can be viewed as the hermitian metric ratio $\frac{|\eta_{\alpha}|_{g}}{|\eta_{\alpha}|_{h}}$. Since $\sqrt{-1}\partial\bar{\partial}$ of the plurisubharmonic function $\varphi_{\alpha}$ is a closed positive current,
$\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha}})=e^{\varphi_{\alpha}}(\sqrt{-1}\partial\varphi_{\alpha}\wedge\bar{\partial}\varphi_{\alpha}+\sqrt{-1}\partial\bar{\partial}\varphi_{\alpha})$
is also a closed positive current on $U_{\alpha}$. This shows that $\partial\varphi_{\alpha}$ is $L^{2}$ with respect to the wight $e^{\varphi_{\alpha}}$ and $e^{\varphi_{\alpha}}\partial\bar{\partial}\varphi_{\alpha}$ has locally finite measure corfficients on $U_{\alpha}$. Hence
$\partial\psi=e^{\varphi_{\alpha}}(\partial\varphi_{\alpha}|\eta_{\alpha}|^{2}_{g}+\partial|\eta_{\alpha}|^{2}_{g})$
is $L^{2}$-integrable on $M$. We will derive the following integral inequality for $\psi$, which is the critical point of the proof of Theorem 1.4.
Proposition 3.1.
The setting is as above, then
$-\int_{M}|\partial\psi|^{2}_{g}\cdot\omega^{n}\geq\frac{p^{2}}{2}\int_{M}\kappa\psi^{2}\cdot\omega^{n}$ (3.1)
where $\kappa$ is a continuous function such that $\kappa(x)$ which is a lower bound of the holomorphic sectional curvature of $\omega$ at $x$.
3.1. Proof of the integral inequality (3.1)
This subsection is devoted to the proof of Proposition 3.1. Firstly, we will proof the following key identity for forms.
Lemma 3.2.
Let $(M,g)$ be a Hermitian manifold of complex dimension $n$, $\alpha$ be a real (1,1)-form and $\eta$ be a (p,0)-form on $M$. We can define a real semi-positive $(1,1)$-form $\beta$ associated to $\eta$ by
$\beta=\operatorname{\Lambda}^{p-1}\left((\sqrt{-1})^{p^{2}}\frac{\eta\wedge\bar{\eta}}{p!}\right),$ (3.2)
where $\operatorname{\Lambda}$ is the dual Lefschetz operator with respect to $\omega$. Then
$\operatorname{tr}_{\omega}\beta=|\eta|^{2}_{g}$ (3.3)
and
$\begin{split}(\sqrt{-1})^{p^{2}}\alpha\wedge\eta\wedge\bar{\eta}\wedge\frac{\omega^{n-p-1}}{(n-p-1)!}=\left[\operatorname{tr}_{\omega}\alpha\cdot|\eta|_{g}^{2}-p\langle\alpha,\beta\rangle_{g}\right]\frac{\omega^{n}}{n!},\end{split}$ (3.4)
where $\omega$ is the fundamental (1,1)-form associated to the Hermitian metric $g$.
Proof.
In local holomorphic coordinates $\{z^{1},\cdots,z^{n}\}$, we have $\alpha=\sqrt{-1}\alpha_{j\bar{k}}dz^{j}\wedge d\bar{z}^{k}$, $\omega=\sqrt{-1}g_{j\bar{k}}dz^{j}\wedge d\bar{z}^{k}$ and $\eta=\eta_{I_{p}}dz^{I_{p}}$, where $I_{p}=(i_{1},\cdots,i_{p})\in\mathbb{N}_{+}^{p}$ and $dz^{I_{p}}=dz^{i_{1}}\wedge\cdots\wedge dz^{i_{p}}$. For any $x_{0}\in M$, we can choose local holomorphic coordinates centered at $x_{0}$ such that
$\omega=\sqrt{-1}\sum\limits_{i}dz^{i}\wedge d\bar{z}^{i}$
and
$\alpha=\sqrt{-1}\sum\limits_{i}\alpha_{i\bar{i}}dz^{i}\wedge d\bar{z}^{i}$
at $x_{0}$. So, we have
$\omega^{n-p-1}$
$=$ $(\sqrt{-1})^{n-p-1}(n-p-1)!\sum\limits_{i_{1}<i_{2}<\cdots<i_{n-p-1}}dz^{i_{1}}\wedge d\bar{z}^{i_{1}}\wedge\cdots\wedge dz^{i_{n-p-1}}\wedge d\bar{z}^{i_{n-p-1}},$
and
$(\sqrt{-1})^{p^{2}}\alpha\wedge\frac{\eta\wedge\bar{\eta}}{p!}$
$=$ $(\sqrt{-1})^{p+1}\sum\limits_{J_{p}}\sum\limits_{i\notin J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}dz^{i}\wedge d\bar{z}^{i}\wedge dz^{j_{1}}\wedge d\bar{z}^{j_{1}}\wedge\cdots\wedge dz^{j_{p}}\wedge d\bar{z}^{j_{p}}$
+(terms involving ηJpηKp¯ with Jpπ(Kp) for some permutation $+(\text{terms involving $\eta_{J_{p}}\overline{\eta_{K_{p}}}$ with $J_{p}\neq\pi(K_{p})$ for some permutation $\pi$}).$
It follows that
$\begin{split}&(\sqrt{-1})^{p^{2}}\frac{n!}{(n-p-1)!}\alpha\wedge\frac{\eta\wedge\bar{\eta}}{p!}\wedge\omega^{n-p-1}\\ =&(\sqrt{-1})^{n}n!\sum\limits_{J_{p}}\sum\limits_{i\notin J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}dz^{1}\wedge d\bar{z}^{1}\wedge\cdots\wedge dz^{n}\wedge d\bar{z}^{n}\\ =&\left(\sum\limits_{J_{p}}\sum\limits_{i\notin J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}\right)\omega^{n}\\ =&\left(\sum\limits_{i,J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}-\sum\limits_{i\in J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}\right)\omega^{n}\\ =&\left(\sum\limits_{i,J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}-p\sum\limits_{i,I_{p-1}}\alpha_{i\bar{i}}\eta_{iI_{p-1}}\overline{\eta_{iI_{p-1}}}\right)\omega^{n}\\ =&\left[\frac{\operatorname{tr}_{\omega}\alpha\cdot|\eta|_{g}^{2}}{p!}-\langle\alpha\wedge\omega^{p-1},(\sqrt{-1})^{p^{2}}\frac{\eta\wedge\bar{\eta}}{p!(p-1)!}\rangle_{g}\right]\omega^{n}\\ =&\left[\frac{\operatorname{tr}_{\omega}\alpha\cdot|\eta|^{2}_{g}}{p!}-\frac{1}{(p-1)!}\langle\alpha,\operatorname{\Lambda}^{p-1}\left((\sqrt{-1})^{p^{2}}\frac{\eta\wedge\bar{\eta}}{p!}\right)\rangle_{g}\right]\omega^{n}\\ =&\left[\frac{\operatorname{tr}_{\omega}\alpha\cdot|\eta|^{2}_{g}}{p!}-\frac{1}{(p-1)!}\langle\alpha,\beta\rangle_{g}\right]\omega^{n}.\end{split}$ (3.5)
This verifies (3.4). Locally, $\beta=\sqrt{-1}\beta_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}$ and
$\beta_{i\bar{j}}=p!\sum\limits_{I_{p-1}}\eta_{iI_{p-1}}\overline{\eta_{jI_{p-1}}}.$ (3.6)
We can reselect the local holomorphic coordiates such that $(g_{i\bar{j}})$ is still the identity and $(\beta_{i\bar{j}})$ is diagonal at $x_{0}$. Then
$\beta_{i\bar{i}}=p!\sum\limits_{I_{p-1}}\eta_{iI_{p-1}}\overline{\eta_{iI_{p-1}}}\geq 0.$ (3.7)
Hence $\beta$ is semi-positive and
$\operatorname{tr}_{\omega}\beta=p!\sum\limits_{i,I_{p-1}}\eta_{iI_{p-1}}\overline{\eta_{iI_{p-1}}}=|\eta|^{2}_{g}.$ (3.8)
The proof is completed. ∎
Proof of Proposition 3.1.
Now let us proceed to show Proposition 3.1. It follows from Lemma 3.2 that
$\begin{split}n&\psi\sqrt{-1}\partial\bar{\partial}\psi\wedge\omega^{n-1}=e^{\varphi_{\alpha}}\operatorname{tr}_{\omega}(\sqrt{-1}\partial\bar{\partial}\psi)|\eta_{\alpha}|^{2}_{g}\cdot\omega^{n}\\ =&p\langle\sqrt{-1}\partial\bar{\partial}\psi,e^{\varphi_{\alpha}}\beta_{\alpha}\rangle_{g}\cdot\omega^{n}+(\sqrt{-1})^{p^{2}}\frac{n!}{(n-p-1)!}\sqrt{-1}\partial\bar{\partial}\psi\wedge e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}}\wedge\omega^{n-p-1}\end{split}$ (3.9)
where $\beta_{\alpha}=\Lambda^{p-1}(\frac{\eta_{\alpha}\wedge\overline{\eta_{\alpha}}}{(p-1)!})$. $e^{\varphi_{\alpha}}\beta_{\alpha}$ and $e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}}$ are globally defined differential forms with bounded measurable cofficients. For clarity, the arguement will be divided into three steps.
Step 1. The integration by parts
Although $\varphi_{\alpha}$ is in general singular, we can still do the integration by parts as in [12]. For reader’s convenience, we give a detailed explaination. Take a partitions of unity $\{\rho_{\alpha}\}$ subordinated to the finite cover $\{U_{\alpha}\}$. Since $\mathrm{supp}\rho_{\alpha}\subset\subset U_{\alpha}$, for any $\epsilon>0$ sufficiently small, there exists a sequence of smooth plurisubharmonic functions $\varphi_{\alpha,\epsilon}$ decreasing to $\varphi_{\alpha}$ on some open subset $V_{\alpha}$ such that $\mathrm{supp}\rho_{\alpha}\subset\subset V_{\alpha}\subset\subset U_{\alpha}$. Now the results of Bedford-Taylor (Lemma 2.4) applied to the uniformly bounded function $e^{\varphi_{\alpha,\epsilon}}$ on $V_{\alpha}$, imply that we have local weak covergence
$e^{\varphi_{\alpha,\epsilon}}\partial\bar{\partial}\varphi_{\alpha,\epsilon}\rightarrow e^{\varphi_{\alpha}}\partial\bar{\partial}\varphi_{\alpha},\ e^{\varphi_{\alpha,\epsilon}}\rightarrow e^{\varphi_{\alpha}}\ \text{and}\ e^{\varphi_{\alpha,\epsilon}}\partial\varphi_{\alpha,\epsilon}\wedge\bar{\partial}\varphi_{\alpha,\epsilon}\rightarrow e^{\varphi_{\alpha}}\partial\varphi_{\alpha}\wedge\bar{\partial}\varphi_{\alpha}$
on $V_{\alpha}$, then
$\begin{split}&-\int_{M}\left|\partial\psi\right|^{2}_{g}\cdot\omega^{n}=-n\int_{M}\partial\psi\wedge\bar{\partial}\psi\wedge\sum\limits_{\alpha}\rho_{\alpha}\omega^{n-1}\\ =&-n\sum\limits_{\alpha}\int_{V_{\alpha}}\sqrt{-1}\partial(e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2})\wedge\bar{\partial}(e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2})\wedge\rho_{\alpha}\omega^{n-1}\\ =&-n\sum\limits_{\alpha}\lim\limits_{\epsilon\rightarrow 0}\int_{V_{\alpha}}\sqrt{-1}\partial(e^{\varphi_{\alpha},\epsilon}|\eta_{\alpha}|_{g}^{2})\wedge\bar{\partial}(e^{\varphi_{\alpha,\epsilon}}|\eta_{\alpha}|_{g}^{2})\wedge\rho_{\alpha}\omega^{n-1}\\ =&n\sum\limits_{\alpha}\lim\limits_{\epsilon\rightarrow 0}\int_{V_{\alpha}}e^{\varphi_{\alpha},\epsilon}|\eta_{\alpha}|_{g}^{2}\left[\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha,\epsilon}}|\eta_{\alpha}|_{g}^{2})\wedge\rho_{\alpha}\omega^{n-1}+\partial\rho_{\alpha}\wedge\sqrt{-1}\bar{\partial}(e^{\varphi_{\alpha,\epsilon}}|\eta_{\alpha}|_{g}^{2})\wedge\omega^{n-1}\right]\\ =&n\sum\limits_{\alpha}\int_{V_{\alpha}}e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2}\left[\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2})\wedge\rho_{\alpha}\omega^{n-1}+\partial\rho_{\alpha}\wedge\sqrt{-1}\bar{\partial}(e^{\varphi_{\alpha}}|\eta_{\alpha}|_{g}^{2})\wedge\omega^{n-1}\right]\\ =&n\int_{M}\psi\sqrt{-1}\partial\bar{\partial}\psi\wedge\omega^{n-1}.\end{split}$ (3.10)
The last equality follows from the fact $\sum\limits_{\alpha}\rho_{\alpha}=1$. Combine with (3.9), we get
$\begin{split}0\geq&-\int_{M}|\partial\psi|_{g}^{2}\cdot\omega^{n}\\ =&p\int_{M}\langle\sqrt{-1}\partial\bar{\partial}\psi,e^{\varphi_{\alpha}}\beta_{\alpha}\rangle_{g}\cdot\omega^{n}+(\sqrt{-1})^{p^{2}}\frac{n!}{(n-p-1)!}\int_{M}\sqrt{-1}\partial\bar{\partial}\psi\wedge e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}}\wedge\omega^{n-p-1}.\end{split}$ (3.11)
Next, let us separately analysis each term on the right-hand side of integral inequality (3.11). Firstly, we consider the sign of first term of (3.11).
Step 2. The Bochner formula.
For any $x_{0}\in U_{\alpha}$, we choose local holomorphic coordinates centered at $x_{0}$ such that
$\omega=\sqrt{-1}\sum\limits_{i}dz^{i}\wedge d\bar{z}^{i}$
and
$\beta_{\alpha}=\sqrt{-1}\sum\limits_{i}\beta_{i\bar{i}}dz^{i}\wedge d\bar{z}^{i}$
at $x_{0}$. For simplicity, we omit the index $\alpha$ in the following calculation. We write $\eta=\sum\limits\eta_{I_{p}}dz^{I_{p}}$, recall (3.3) and (3.7) that we have
$\operatorname{tr}_{\omega}\beta=|\eta|_{g}^{2}$ (3.12)
and
$\beta_{i\bar{i}}=p!\sum\limits_{I_{p-1}}\eta_{iI_{p-1}}\overline{\eta_{iI_{p-1}}}\geq 0.$ (3.13)
By a direct calculation, we derive the following Bochner type formula in the sense of distribution,
$\begin{split}\partial_{u}\partial_{\bar{u}}(e^{\varphi}|\eta|^{2}_{g})=&e^{\varphi}(|\eta|^{2}\partial_{u}\partial_{\bar{u}}\varphi+|D^{\prime}_{u}\eta+\eta\partial_{u}\varphi|_{\Lambda^{p,0}M}^{2}+p!\sum\limits_{I_{p}}\sum\limits_{k=1}^{p}\sum\limits_{l=1}^{n}R_{u\bar{u}i_{k}\bar{l}}\eta_{I_{p}}\overline{\eta_{i_{1}\cdots(l)_{k}\cdots i_{p}}})\\ =&e^{\varphi}(|\eta|^{2}\partial_{u}\partial_{\bar{u}}\varphi+|D^{\prime}_{u}\eta+\eta\partial_{u}\varphi|_{\Lambda^{p,0}M}^{2}+p(p!)\sum\limits_{l,k=1}^{n}R_{u\bar{u}l\bar{k}}\sum\limits_{J_{p-1}}\eta_{lJ_{p-1}}\overline{\eta_{kJ_{p-1}}})\\ =&e^{\varphi}(|\eta|^{2}\partial_{u}\partial_{\bar{u}}\varphi+|D^{\prime}_{u}\eta+\eta\partial_{u}\varphi|_{\Lambda^{p,0}M}^{2}+p\sum\limits_{l,k}R_{u\bar{u}l\bar{k}}\beta_{l\bar{k}})\end{split}$ (3.14)
for all $u\in T^{1,0}_{x_{0}}M$, where $R_{i\bar{j}k\bar{l}}=R(\partial_{i},\overline{\partial_{j}},\partial_{k},\overline{\partial_{l}})$ and $D^{\prime}$ is the $(1,0)$-part of the connection $D$ of $\Lambda^{p,0}M$ inducced by Chern connection of $(T^{1,0}M,\omega)$. The second equality follows from interchanging the order of summation. Hence
$\begin{split}&p\left\langle\sqrt{-1}\partial\bar{\partial}\psi,e^{\varphi}\beta\right\rangle_{g}=e^{2\varphi}(p|\eta|^{2}\partial_{i}\partial_{\bar{i}}\varphi\beta_{i\bar{i}}+p\sum\limits_{i}|D^{\prime}_{i}\eta+\eta\partial_{i}\varphi|^{2}\beta_{i\bar{i}}+p^{2}\sum\limits_{i,k}R_{i\bar{i}k\bar{k}}\beta_{i\bar{i}}\beta_{k\bar{k}})\\ =&e^{2\varphi}\left(p\psi\langle\sqrt{-1}\partial\bar{\partial}\varphi,\beta\rangle+p\left<\sqrt{-1}\langle D^{\prime}\eta+\eta\partial\varphi,\overline{D^{\prime}\eta+\eta\partial\varphi}\rangle_{g},\beta\right>_{g}+p^{2}\langle\operatorname{Rm},-\beta\otimes\beta\rangle_{g}\right).\end{split}$ (3.15)
It follows from Royden’s observation(Lemma 2.6) that
$\begin{split}e^{2\varphi}\langle\operatorname{Rm},-\beta\otimes\beta\rangle_{g}=&e^{2\varphi}\sum\limits_{i,k}R_{i\bar{i}k\bar{k}}\beta_{i\bar{i}}\beta_{k\bar{k}}\\ \geq&e^{2\varphi}\frac{\kappa(x_{0})}{2}\left[(\sum\limits_{i}\beta_{i\bar{i}})^{2}+\sum\limits_{i}\beta_{i\bar{i}}^{2}\right]\\ =&\frac{\kappa(x_{0})}{2}e^{2\varphi}\left(|\eta|_{g}^{4}+|\beta|_{g}^{2}\right)\\ \geq&\frac{\kappa(x_{0})}{2}\psi^{2}\end{split}$ (3.16)
at $x_{0}$. The fact that $\sqrt{-1}\partial\bar{\partial}\varphi\geq 0$ and $\beta\geq 0$ implies that
$e^{2\varphi}(p\psi\langle\partial\bar{\partial}\varphi,\beta\rangle+p\left<\sqrt{-1}\langle D^{\prime}\eta+\eta\partial\varphi,\overline{D^{\prime}\eta+\eta\partial\varphi}\rangle_{g},\beta\right>_{g}\geq 0.$ (3.17)
Hence, we have
$p\int_{M}\left\langle\sqrt{-1}\partial\bar{\partial}\psi,e^{\varphi}\beta\right\rangle_{g}\cdot\omega^{n}=p\int_{M}\left\langle\sqrt{-1}\partial\bar{\partial}\psi,e^{\varphi}\beta\right\rangle_{g}\cdot\omega^{n}\geq\frac{p^{2}}{2}\int_{M}\kappa\psi^{2}\cdot\omega^{n}.$
Step 3. The sign of the second term of (3.11):
As
¯$\partial\bar{\partial}(e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}})=\partial(e^{\varphi_{\alpha}}\bar{\partial}\varphi_{\alpha}\wedge\eta_{\alpha}\wedge\overline{\eta_{\alpha}}+(-1)^{p}\eta_{\alpha}\wedge\overline{\partial\eta_{\alpha}})$
$=$ $(\sqrt{-1})^{2p}e^{\varphi_{\alpha}}(\partial\eta_{\alpha}+\partial\varphi_{\alpha}\wedge\eta_{\alpha})\wedge\overline{(\partial\eta_{\alpha}+\partial\varphi_{\alpha}\wedge\eta_{\alpha})}+e^{\varphi_{\alpha}}\partial\bar{\partial}\varphi_{\alpha}\wedge\eta_{\alpha}\wedge\overline{\eta_{\alpha}},$
by applying the integration by parts which is valid as previously explained, we get
$(\sqrt{-1})^{p^{2}}\int_{M}\sqrt{-1}\partial\bar{\partial}\psi\wedge e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\bar{\eta}_{\alpha}\wedge\omega^{n-p-1}$
$=$ $(\sqrt{-1})^{p^{2}}\int_{M}\psi\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta}_{\alpha}\wedge\omega^{n-p-1})$
$=$ $(\sqrt{-1})^{p^{2}}\int_{M}\psi\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta}_{\alpha})\wedge\omega^{n-p-1}$
$=$ $\int_{M}\psi(\sqrt{-1})^{(p+1)^{2}}e^{\varphi_{\alpha}}\gamma_{\alpha}\wedge\overline{\gamma_{\alpha}}\wedge\omega^{n-p-1}+\int_{M}\psi\sqrt{-1}\partial\bar{\partial}\varphi_{\alpha}\wedge e^{\varphi_{\alpha}}(\sqrt{-1})^{p^{2}}\eta_{\alpha}\wedge\bar{\eta}_{\alpha}\wedge\omega^{n-p-1}$
$\geq$ $0$
where $\gamma_{\alpha}=\partial\eta_{\alpha}+\partial\varphi_{\alpha}\wedge\eta_{\alpha}$. The last inequality follows from the fact
$\sqrt{-1}\partial\bar{\partial}\varphi_{\alpha}\geq 0,\ (\sqrt{-1})^{(p+1)^{2}}\gamma_{\alpha}\wedge\overline{\gamma_{\alpha}}\geq 0\text{ and }(\sqrt{-1})^{p^{2}}\eta_{\alpha}\wedge\bar{\eta}_{\alpha}\geq 0.$
In summary, we get the desired integral inequality. The proof is completed. ∎
3.2. The proof of Theorem 1.4
In this subsection, we proceed to show Theorem 1.4. The method is based on performing a pointwise computation.
Proof of Theorem 1.4.
Let $\mathcal{F}$ be an invertible subsheaf of $\mathcal{O}({\Lambda^{p,0}M})$ that is pseudo-effective. Since the holomorphic sectional curvature of $\omega$ is nonnegative on $M$, we can choose $\kappa\equiv 0$ in the setting 3. By Proposition 3.1, we have
$-\int_{M}|\partial\psi|_{g}^{2}\cdot\omega^{n}\geq 0.$
Hence $d\psi=0$ almost everywhere on $M$, which implies that $\psi=C$ for some constant $C\geq 0$ almost everywhere. We claim that $C>0$. Suppose not, we have $\psi=0$ almost everywhere on $U_{\alpha}$. Since $\varphi_{\alpha}\in Psh(U_{\alpha})\cap L_{loc}^{1}(U_{\alpha})$, $\{\varphi_{\alpha}=-\infty\}$ has zero Lebesgue measure. This fact implies that $\eta_{\alpha}=0$ almost everywhere on $U_{\alpha}$. Then $\eta_{\alpha}\equiv 0$ on $U_{\alpha}$ due to the smoothness of $\eta_{\alpha}$, which contradicts that $\eta_{\alpha}$ generates $\mathcal{F}|_{U_{\alpha}}$. Therefore, $\varphi_{\alpha}<+\infty$ and $e^{\varphi_{\alpha}}|\eta_{\alpha}|^{2}_{g}=C>0$ implies that $|\eta_{\alpha}|^{2}_{g}>0$ and $\varphi_{\alpha}=\log C-\log|\eta_{\alpha}|^{2}_{g}$ is a smooth function on $U_{\alpha}$. The following calculation will be done in the smooth sense.
For any $\alpha$, take an arbitrary point $x_{0}$ of $U_{\alpha}$. For simplicity, we omit the index $\alpha$. Recall that we can define a real semi-positive $(1,1)$-form $\beta$ associated to $\eta$ by
$\beta=\operatorname{\Lambda}^{p-1}\left((\sqrt{-1})^{p^{2}}\frac{\eta\wedge\overline{\eta}}{p!}\right)$ (3.18)
such that
$\operatorname{tr}_{\omega}\beta=|\eta|^{2}_{g}>0.$ (3.19)
We choose local holomorphic coordinates centered at $x_{0}$ such that $(g_{i\bar{j}})$ is the identity and $(\beta_{i\bar{j}})$ is diagonal at $x_{0}$. Since $\beta_{i\bar{i}}$ are nonnegative for all $1\leq i\leq n$ and $\sum\limits_{i}\beta_{i\bar{i}}>0$, without loss of generality, we may assume that
$\beta_{1\bar{1}}\geq\beta_{2\bar{2}}\geq\cdots\geq\beta_{m\bar{m}}>0=\beta_{m+1\overline{m+1}}=\cdots=\beta_{n\bar{n}}$ (3.20)
for some positive integer $m\leq n$. Denote
[…]
From the Bochner type formula (3.14), we have
$0=e^{-\varphi}\partial_{v}\partial_{\bar{v}}(e^{\varphi}|\eta|^{2}_{g})=|\eta|^{2}_{g}\partial_{v}\partial_{\bar{v}}\varphi+|D^{\prime}_{v}\eta+\eta\partial_{v}\varphi|^{2}+p\sum\limits_{j=1}^{m}R_{v\bar{v}j\bar{j}}\beta_{j\bar{j}}\geq p\sum\limits_{j=1}^{m}R_{v\bar{v}j\bar{j}}\beta_{j\bar{j}}\lambda_{j}^{2}$ (3.21)
for all $v\in T_{x_{0}}^{1,0}M$. Let $t=(t_{1},\cdots,t_{m})$ be the complex coordinates of $\mathbb{C}^{m}$ and $d\theta$ be the standard volume measure of the unit sphere $S_{2m-1}:=\{t\in\mathbb{C}^{m}:\sum\limits_{i=1}^{m}|t_{i}|^{2}=1\}$. By the Berger’s averaging trick ([2]), we get
[…]
For any $t\in S_{2m-1}$,
$\sum\limits_{i,j,k,l=1}^{m}R_{i\bar{j}k\bar{l}}\lambda_{i}t_{i}\overline{\lambda_{j}}\overline{t_{j}}\lambda_{k}t_{k}\overline{\lambda_{l}}\overline{t_{l}}=R(V_{t},\overline{V_{t}},V_{t},\overline{V_{t}})\geq 0,$ (3.23)
where $V_{t}:=\sum\limits_{i=1}^{m}\lambda_{i}t_{i}\frac{\partial}{\partial z^{i}}$. Combining (3.22) with (3.23), $R(V_{t},\overline{V_{t}},V_{t},\overline{V_{t}})=0,\ \forall t\in S_{2m-1}$. Taking $t_{0}\in S_{2m-1}$ such that only the $l$-th component is $1$ and all other components are 0, then $V_{t_{0}}=\lambda_{l}\frac{\partial}{\partial z^{l}}$ and
$R_{l\bar{l}l\bar{l}}=\frac{1}{\lambda_{l}^{4}}R(V_{t_{0}},\overline{V_{t_{0}}},V_{t_{0}},\overline{V_{t_{0}}})=0,\ \forall 1\leq l\leq m.$ (3.24)
Since the holomorphic sectional curvature is nonnegative, $\frac{\partial}{\partial z^{l}}$ minimizes the holomorphic sectional curvature of $\omega$ at $x_{0}$. It follows from Lemma 2.7 that
$R_{l\bar{l}v\bar{v}}\geq 0,\ \forall 1\leq l\leq m,v\in T^{1,0}_{x_{0}}M.$
Since $\partial_{v}\partial_{\bar{v}}\varphi\geq 0$, (3.21) implies that $R_{l\bar{l}v\bar{v}}=0,\ \forall 1\leq l\leq m$ and
$\partial_{v}\partial_{\bar{v}}\varphi=0,\ D^{\prime}_{v}\eta+\eta\partial_{v}\varphi=0$
[…]
Since $x_{0}\in M$ is arbitrary, we have ¯$\partial\bar{\partial}\varphi=0$ on $M$ and $D^{\prime}\eta+\eta\partial\varphi=0$ on $M$. ¯$\partial\bar{\partial}\varphi_{\alpha}=0$ implies that $(\mathcal{F},e^{-\varphi})$ is a flat line bundle. In each coordinate neighborhood $U_{\alpha}$, $\varphi_{\alpha}$ is pluriharmonic, then $\varphi_{\alpha}$ can be written $\varphi_{\alpha}=f_{\alpha}+\overline{f_{\alpha}}$ for some holomorphic function $f_{\alpha}\in\mathcal{O}(U_{\alpha})$ and
$D^{\prime}(e^{f_{\alpha}}\eta_{\alpha})=e^{f_{\alpha}}(D^{\prime}\eta_{\alpha}+\eta_{\alpha}\partial\varphi_{\alpha})=0$
where $e^{f_{\alpha}}\eta_{\alpha}$ is a local holomorphic section of $\mathcal{F}|_{U_{\alpha}}$. This shows that $\mathcal{F}$ is invariant by parallel transport induced by connection $D$ of $\Lambda^{p,0}M$. The proof is completed. ∎
4. Proof of Theorem 1.5 and Theorem 1.7
In this subsection, we proceed to show Theorem 1.5 and Theorem 1.7. The proof of Theorem 1.5 is based on Theorem 1.4 (1) and Campana-Demailly-Peternell’s critetion. The proof of Theorem 1.7 follows the arguement in [7, section 4] for $\operatorname{Ric}\geq 0$, demonstrating that factors of irreducible holonomy $H=U(n)$ are rationally connected manifolds. The difference is that factors of irreducible holonomy $H=SU(n)$ or $Sp(\frac{n}{2})$ can be excluded by statement (1) of Theorem 1.4.
Proof of Theorem 1.5.
Since the holomorphic sectional curvature is nonnegative and there exists no nonzero truly flat tangent vector, any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ is not pseudo-effective by the statement (1) of Theorem 1.4. This fact implies that $H^{p,0}_{\bar{\partial}}(M)=0,\ \forall 1\leq p\leq n$, in particular, M is projective by Kodaira’s theorem ([21]). Combine with Campana-Demailly-Peternell’s criterion (Lemma 2.5), $M$ is rationally connected. ∎
Proof of Theorem 1.7.
Let $(M,\omega)\simeq\prod(M_{l},\omega_{l})$ be the De Rham decomposition of $(M,\omega)$, induced by a decomposition of the holonomy representation in irreducible representations ([4]). Since the holonomy is contained in $U(n)$, all factors $(M_{l},\omega_{l})$ are Kähler manifolds with irreducible holonomy and holonomy group $H_{l}\subset U(n_{l})$, $n_{l}=\dim M_{l}$. The product structure implies that $\omega_{l}$ has nonnegative holomorphic sectional curvature. By Berger’s classification of holonomy groups ([4]), there are only three possibilities: $H_{l}=U(n_{l}),\ SU(n_{l})$ or $Sp(\frac{n_{l}}{2})$, unless $M_{l}$ is a Hermitian symmetric space, then necessarily of compact type (such symmetric space were classified by Élie Cartan, they are rational homogeneous, hence rationally connected, their holonomy groups are also known, see. e.g. [5]).
Case 1: $H_{l}=SU(n_{l})$ or $Sp(\frac{n_{l}}{2})$. Then $M_{l}$ is a Calabi-Yau manifold or a holomorphic symplectic manifold (see e.g. [1]). Recall that a simply connected Calabi-Yau manifold $M_{l}$ is a compact simply connected projective manifold with $K_{M_{l}}\simeq O_{M_{l}}$ and $H^{0}(M_{l},\Omega_{M_{l}}^{p})=0$ for $0<p<\dim M_{l}$. A holomorphic symplectic manifold is a compact Kähler manifold $X$ which admits a holomorphic symplectic $(2,0)$-form (of maximal rank everywhere). In particular, $M_{l}$ admits a holomorphic $(n_{l},0)$-form that vanishes nowhere. Recall that we can define a nonnegative $(1,1)$-form $\beta=\Lambda^{p-1}({\frac{\eta\wedge\overline{\eta}}{n_{l}!}})$. For any point $x_{0}$ of $M_{l}$, we can choose some local holomorphic coordinates such that the coefficients of $w_{l}$ is identity and $\beta$ is diagonal at $x_{0}$. Locally, we can write $\eta=\frac{1}{n_{l}!}\sum\limits fdz^{1}\wedge\cdots dz^{n}$. It follows from the local representation (3.7) of $\beta$ that we have
$\beta=\frac{|\eta|^{2}}{n_{l}}\sum\limits_{i=1}^{n_{l}}dz^{i}\wedge d\overline{z}^{i}$
at $x_{0}$. The proof of Theorem 1.4 tells us that $\{\frac{\partial}{\partial z_{i}}\}_{i=1}^{n_{l}}$ are truly flat tangent vectors at $x_{0}$. Hence the Riemannian curvature tensor $\operatorname{Rm}$ of $\omega_{l}$ vanishes at $x_{0}$. Since $x_{0}$ is arbitrary, $\operatorname{Rm}$ vanishes everywhere, $M_{l}$ is a simply connected complete manifold with zero holomorphic sectional curvature. By classification of complex space forms, $M_{l}$ must be holomorphically isometric to complex Euclidean space $\mathbb{C}^{n_{l}}$, which contradicts the fact that $M_{l}$ is compact.
Case 2: $H_{l}=U(n_{l})$. Since the natural representation of $U(n_{l})$ on $(\mathbb{C}^{n_{l}})^{\otimes p}$ has no invariant line on which $U(n_{l})$ induces a trivial action, there cannot exist any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M_{i})$, $1\leq p\leq n$ such that $\mathcal{F}$ is invariant by parallel transport induced by the connection $D$. Combine with statement (2) of Theorem 1.4, any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ is not pseudo-effective. Therefore $M_{l}$ is rationally connected by Campana-Demailly-Peternell’s criterion.
In summary, each $(M_{l},\omega_{l})$ is a rationally connected manifold with nonnegative holomorphic sectional curvature and irreducible holonomy (holonomy $U(n_{l})$ unless $M_{l}$ is a hermitian symmetric space of compact type). ∎
5. Proof of Theorem 1.10
In this section, we give a proof of Theorem 1.10.
Proof of Theorem 1.10.
Denote $n=\dim M$ and $r=\operatorname{rank}E$. For any $t\in\mathbb{N}_{+},\ m\in[0,k]\bigcap\mathbb{N}$ and $I_{k}=(i_{1},\cdots,i_{k})\in\mathbb{N}_{+}^{k}$ for some $k\in\mathbb{N}$, we define
$I_{k}(m,t)=(i_{1},\cdots,i_{m-1},t,i_{m},\cdots,i_{k})\in\mathbb{N}_{+}^{k+1}.$
We equip $E$ with a Hermitian metric $H$ and denote its chern curvature tensor by $R^{H}$. Set
$\delta_{2}=\min\limits_{x\in M}\min_{u\in T^{1,0}_{x}M\setminus 0,V\in E_{x}\setminus 0}-\frac{R^{H}(u,\bar{u},V,\bar{V})}{|u|^{2}_{h}|V|^{2}_{H}}.$ (5.1)
Since $(M,h)$ has positive real bisectional curvature, there exists $\delta_{1}>0$ such that for any $x\in M$, any unitary basis $\{\frac{\partial}{\partial z^{1}},\cdots,\frac{\partial}{\partial z^{n}}\}$ of $T_{x}^{1.0}M$ and any nonnegative constants $a=\{a_{1},\cdots,a_{n}\}$ with $|a|^{2}=a_{1}^{2}+\cdots+a_{n}^{2}>0$, we have that
$\sum\limits_{i,j=1}^{n}R_{i\bar{i}j\bar{j}}a_{i}a_{j}>\delta_{1}|a|^{2}.$ (5.2)
For any $x\in M$, we choose local holomorphic coordinates $(z^{1},\cdots,z^{n})$ centered at $x$ such that $h_{i\bar{j}}$ is the identity at $x$. Let $\psi$ be any holomorphic section of $H^{0}(M,((T^{1,0}M)^{*})^{\otimes p}\otimes E^{\otimes l})$. Locally
$\psi=\sum\limits_{I_{p},J_{l}}a_{I_{p},J_{l}}dz^{i_{1}}\otimes\cdots\otimes dz^{i_{p}}\otimes e^{j_{1}}\otimes\cdots\otimes e^{j_{l}}$
with $I_{p}=(i_{1},\cdots,i_{p})\in\mathbb{N}_{+}^{p},J_{l}=(j_{1},\cdots,j_{l})\in\mathbb{N}_{+}^{l}$, where $\{e^{k}\}_{k=1}^{r}$ is a local holomorphic basis of $E$ such that $(H_{t\bar{s}})$ is diagonal at $x$. By a direct computation, we have the following Bochner type formula,
$\partial_{u}\partial_{\bar{u}}|\psi|^{2}$
$=$ $\langle D^{\prime}_{u}\psi,\overline{D^{\prime}_{u}\psi}\rangle+\sum\limits_{i,j=1}^{n}R^{h}_{u\bar{u}i\bar{j}}\sum\limits_{m=1}^{p}\sum\limits_{I_{p-1},J_{l}}a_{I_{p-1}(m,i),J_{l}}\overline{a_{I_{p-1}(m,j),J_{l}}}$
$-\sum\limits_{t,s=1}^{r}R^{H}_{u\bar{u}t\bar{s}}\sum\limits_{k=1}^{l}\sum\limits_{I_{p},J_{l-1}}a_{I_{p},J_{l-1}(k,t)}\overline{a_{I_{p},J_{l-1}(k,s)}}$
holds for all $u\in T_{x}^{1,0}M$. Define a real nonnegative $(1,1)$-form $\alpha=\sqrt{-1}\alpha_{i\bar{j}}dz^{i}\wedge d\bar{z}^{j}$ by
$\alpha_{i\bar{j}}=\sum\limits_{m=1}^{p}\sum\limits_{I_{p-1},J_{l}}a_{I_{p-1}(m,i),J_{l}}\overline{a_{I_{p-1}(m,j),J_{l}}}=\sum\limits_{m=1}^{p}\langle\iota^{m}_{\frac{\partial}{\partial z^{i}}}(\psi),\overline{\iota^{m}_{\frac{\partial}{\partial z^{j}}}(\psi)}\rangle$ (5.3)
with $\iota^{m}_{\frac{\partial}{\partial z^{i}}}(\psi)=\psi(\cdots,\frac{\partial}{\partial z^{i}},\cdots;\cdots)$ where the contractions are done at the $m$-th slot. In a similar way, define a section $\beta=\beta_{t\bar{s}}e^{t}\otimes\overline{e^{s}}$ of $E\otimes\overline{E}$ such that the components $(\beta_{t\bar{s}})$ is a nonnegative Hermitian matrix by
$\beta_{t\bar{s}}=\sum\limits_{k=1}^{l}\sum\limits_{I_{p},J_{l-1}}a_{I_{p},J_{l-1}(k,t)}\overline{a_{I_{p},J_{l-1}(k,s)}}=\sum\limits_{k=1}^{l}\langle\iota^{p+k}_{(e^{t})^{*}}(\psi),\overline{\iota^{p+k}_{(e^{s})^{*}}(\psi)}\rangle$ (5.4)
with $\iota^{p+k}_{(e^{t})^{*}}(\psi)=\psi(\cdots;\cdots,(e^{t})^{*},\cdots)$ where the contractions are done at the $(p+k)$-th slot. Now we can reselect the local holomorphic coordinates of $M$ and the local holomorphic basis of $E$ such that $(h_{i\bar{j}}),(H_{t\bar{s}})$ are still the identity and $(\alpha_{i\bar{j}}),(\beta_{t\bar{s}})$ are diagonal at $x$, then
$\sum\limits_{i=1}^{n}\alpha_{i\bar{i}}=\sum\limits_{i=1}^{n}\sum\limits_{m=1}^{p}\sum\limits_{I_{p-1},J_{l}}a_{I_{p-1}(m,i),J_{l}}\overline{a_{I_{p-1}(m,i),J_{l}}}=p|\psi|^{2}$ (5.5)
and
$\sum\limits_{t=1}^{r}\beta_{t\bar{t}}=\sum\limits_{t=1}^{r}\sum\limits_{k=1}^{l}\sum\limits_{I_{p},J_{l-1}}a_{I_{p},J_{l-1}(k,t)}\overline{a_{I_{p},J_{l-1}(k,t)}}=l|\psi|^{2}.$ (5.6)
So, for any holomorphic section $\psi\in H^{0}(M,((T^{1,0}M)^{*})^{\otimes p}\otimes E^{\otimes l})$,
$\langle\sqrt{-1}\partial\bar{\partial}|\psi|^{2},\alpha\rangle_{h}=$ $\sum\limits_{i=1}^{n}\partial_{i}\partial_{\bar{i}}|\psi|^{2}\alpha_{i\bar{i}}$
$\geq$ $\sum\limits_{i,j=1}^{n}R^{h}_{i\bar{i}j\bar{j}}\alpha_{i\bar{i}}\alpha_{j\bar{j}}-\sum\limits_{i=1}^{n}\sum\limits_{t=1}^{r}R^{H}_{i\bar{i}t\bar{t}}\alpha_{i\bar{i}}\beta_{t\bar{t}}$
$\geq$ $\delta_{1}\sum\limits_{i=1}^{n}|\alpha_{i\bar{i}}|^{2}+\delta_{2}\sum\limits_{i=1}^{n}\alpha_{i\bar{i}}\sum\limits_{t=1}^{r}\beta_{t\bar{t}}$
$\geq$ $\delta_{1}\frac{(\sum\limits_{i=1}^{n}\alpha_{i\bar{i}})^{2}}{n}+\delta_{2}(\sum\limits_{i=1}^{n}\alpha_{i\bar{i}})(\sum\limits_{t=1}^{r}\beta_{t\bar{t}})$
$=$ $(p+\frac{n\delta_{2}}{\delta_{1}}l)\frac{p\delta_{1}|\psi|^{4}}{n}$
holds for any $x\in M$. Let $C_{E}=\max\{1,-\frac{n\delta_{2}}{\delta_{1}}+2\}$. We assume that $|\psi|^{2}$ achieves its maximum at some point $x_{0}\in M$. If $p,l\in\mathbb{N}_{+}$ with $p\geq C_{E}l$ and $\psi\neq 0$, then we have
$00\geq\langle\sqrt{-1}\partial\bar{\partial}|\psi|^{2},\alpha\rangle_{h}\geq(p+\frac{n\delta_{2}}{\delta_{1}})|\psi|^{4}\geq 2l|\psi|^{4}>$ 0
at $x_{0}$, which is a contradiction. Hence
$H^{0}(M,((T^{1,0}M)^{*})^{\otimes p}\otimes E^{\otimes l})=0\ \text{ for any }p\geq C_{E}l.$ (5.7)
The vanishing theorem (5.7) implies that $H^{2,0}_{\bar{\partial}}(M)=0$. In particular, if $M$ is Kählerian, $M$ is projective by Kodaira’s theorem ([21]). Then (5.7) and Campana-Peternell-Demailly’s criterion (Lemma 2.5) imply that $M$ is rationally connected. The proof is completed. ∎